OpenStudy (anonymous):

Please help...Use the quadratic and a calculator to solve the following equation. Round intermediate calculations and final answers to two decimal places. 6.25y^2 + 12y + 5.76 = 0 i got y = .96, -96 but it was wrong should i round up to 1, -1

OpenStudy (anonymous):

if you require the answer correct to dec places to do not iround intermediate results to 2 dp's

OpenStudy (anonymous):

right so I dont understand the statement Round intermediate calculations and final answers to two decimal places

OpenStudy (anonymous):

2 dec places i meant to say

OpenStudy (anonymous):

Is it bcuz it is saying Use the quadratic and a calculator to solve

OpenStudy (anonymous):

i make the solutios -0.96

OpenStudy (anonymous):

there is only one solution

OpenStudy (anonymous):

so you think incluce the 0

OpenStudy (anonymous):

can a solution set have only one thing in it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

6.25y^2 + 12y + 5.76 = 0 Rationalize the coefficients. \[\frac{25}{4}y^2+12y+\frac{144}{25}=0\]Multiply each side by 100 and simplify.\[625 y^2+1200 y+576=0 \]Use the binomial theorem or, to pull a rabbit out of the hat:\[625 y^2+1200 y+576=(25 y+24)^2 \]\[y=\frac{24}{25}\text{ }\text{or}\text{ }0.96 \]

OpenStudy (anonymous):

if you graph this function it will just touch the x axis at -0.96

OpenStudy (anonymous):

I stand corrected. Yes, the sign is negative. Thank you.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so you are saying that the only solution in the solution set is -0.96??

OpenStudy (anonymous):

There are two solutions, each with identical values.

OpenStudy (anonymous):

so both would be -0.96

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but I probably only need to enter it once?

OpenStudy (anonymous):

the books always say two identical solutions but to be honest i dont understand why

OpenStudy (anonymous):

Woohoo it was correct...thank you both sooooo much!!!

OpenStudy (anonymous):

I use Mathematica to do the calculations.\[\text{Solve}\left[(25 y+24)^2\text{==}0,y\right]=\left\{\left\{y\to -\frac{24}{25}\right\},\left\{y\to -\frac{24}{25}\right\}\right\} \]Notice that it returned two values for y. I would put down as an answer, x = -0.96 twice. Does this answer you questions?

OpenStudy (anonymous):

no probs popcorn

OpenStudy (anonymous):

Your welcome.